# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v3

@@ IV-B Agent Communication Definition
To tackle these conflicts, agent communication is urgently demanded. Specifically, agents need to collaborate with a series of external entities to finish user tasks. In this paper, we present a clear definition of agent communication as follows:
@@
IV-C Agent Communication Classification
Based on the object of communication, we divide agent communication into three classes: user-agent, agent-agent, agent-environment. We will use Figure 4 as a typical example to systematically overview the complete lifecycle of agent communication.
@@ IV-C3 Agent-Environment Communication
Advantages of this classification method. Different entities have essentially differentiated capability characteristics and attack surface attributes. For example, one of the major security risks in user-agent interaction lies in the natural uncontrollability of user input, which is essentially different from agent-agent or agent-environment communication. As a result, classifying agent communication by entity types can directly cluster major vulnerability types and defense strategies that have similar characteristics, providing a structured analysis paradigm for future security research.
IV-D Organization of the Following Sections
@@
As shown in Figure 5, in the following sections, we will discuss agent communication and related security.
In Section V, we will introduce user-agent interaction. Specifically, the risks in this process are divided into two categories: risks from malicious users against benign agents (Section V-B) and risks from compromised agents against benign users (Section V-C). The possible defense countermeasures against malicious users are discussed in Section V-D, while those against malicious agents are discussed uniformly in Section VI-C.
In Section VI, we will classify existing protocols for agent-agent communication. Then, we discuss the risks in Section VI-B and defenses in Section VI-C. Please note that the proposed defenses against compromised agents in this section are also applicable to the risks in Section V-C and Section VII-C.
In Section VII, we first show the related protocols in Section VII-A. Then, the risks in this process are also classified into two categories: risks from malicious environments against benign agents (Section VII-B) and risks from compromised agents against benign environments (Section VII-C). The defenses for compromised agents can be referred to Section VI-C. As a result, we only discuss the defenses for the malicious environments in Section VII-D.
Please note that the risks/defenses we focus on are specific for agent communication instead of LLMs. This is because agents and agent communication still have significant differences from LLMs (as we have discussed in Section III). For example, in the agent collaboration process, the accumulation of tiny benign deviations on each step may lead to an intolerable risk in the final result. Miehling et al. [204] also point out that over-focusing on the capabilities of a single model can result in neglecting the emergent behaviors at the system level and underestimating the true risks. As a result, for LLM-targeted attacks that have valuable inspiration for agent security, we will discuss the typical representatives of them instead of listing all papers exhaustively.
IV-E Takeaways
Conforming to the usage trends of users, the collaboration of multiple agents has become a clear development direction. In this context, agent communication becomes the foundation of the future AI ecosystem. Based on the communication entity, we classify agent communication into three different types: user-agent interaction, agent-agent communication, and agent-environment communication. This classification can naturally distinguish communication with similar vulnerability characteristics, providing a structured research paradigm for future studies.
@@ V User-Agent Interaction
In this section, we will introduce the current user-agent interaction protocols, their security risks, and future defense strategies.
V-A Protocols
PXP. PXP protocol [256] focuses on building an interactive system between human experts and agents in data analysis tasks, targeting issues in complex scientific, medical, and other fields. It is worth mentioning that PXP is not customized for LLM-driven agents, but we think its design has inspirational meaning for agent communication. Therefore, we finally discuss it in this paper. PXP deploys a “two-way intelligibility” mechanism as its core and uses four message tags, namely RATIFY, REFUTE, REVISE, and REJECT, to regulate the interaction between human experts and agents. At the beginning of the interaction, the agent initiates a prediction and provides an explanation first. Subsequently, the two parties communicate alternately. A finite-state machine is used to calculate the message tags and update the context based on the prediction matching (MATCH) and explanation agreement (AGREE) situations. PXP uses a blackboard system to store data, messages, and context information. The process continues until the message limit is reached or specific termination conditions occur. The effectiveness of PXP has been verified in the scenarios of radiology and drug discovery.
Spatial Population Protocols. The Spatial Population Protocols is a minimalist and computationally efficient distributed computing model, specifically designed to solve the Distributed Localization Problem (DLP) in robot systems. Similar to PXP, strictly speaking, this work is not designed for LLM-driven agent systems. However, since it may benefit agents requiring location services, we also discuss it in this paper. Spatial Population Protocols allow agents to obtain pairwise distances or relative position vectors when interacting in Euclidean space. Each agent can store a fixed number of coordinates. During interaction, in addition to exchanging knowledge, geometric queries can also be performed. Through the multi-contact epidemic mechanism, leader election, and self-stabilizing design, it enables n anonymous robots to achieve efficient localization from their respective inconsistent coordinate systems to a unified coordinate consensus, providing a scalable framework for robot collaboration in dynamic environments.
@@ V-A Protocols
AG-UI. AG-UI [217] realizes the communication between users (front-end applications) and agents based on the client-server architecture and completes the communication process by adopting the event-driven mechanism. As shown in Figure 6, the front-end application connects to agents through the AG-UI client (such as a common communication client that supports server-sent events or binary protocols). The client invokes the RUN interface of the protocol layer to send requests to the agent. When the agent processes the request, it generates a streaming event and returns it to the AG-UI client. Event types include lifecycle events (such as start of run, completion of run), text message events (transmitted in segments by start, content, and end), tool call events (passed in the order of start, parameters, and end), and state management events. The AG-UI client handles different types of responses by subscribing to the event stream. Agents can transfer context between each other to maintain the continuity of the conversation. All events follow a unified basic event structure and undergo strict type verification to ensure the reliability and efficiency of communication.
Besides, please note that the previous survey [318] regards CrowdES [24] as an agent-related protocol. After our careful study, we think CrowdES is a framework for generating and evaluating simulated crowds and real-world crowds, which is not suitable for discussion in the field of agent communication. Therefore, we will not list it in this paper.
V-B Security Risk Analysis: Malicious Users Against Benign Agents
@@
| Stage | Risk Type | Representative Threats | Attack Characteristics |
| Malicious users against benign agents | Text-Based Attacks | Prompt Injection | Manipulate agent behavior by malicious prompts, not modifying model |
| Jailbreak Prompts | Bypass security constraints via role-play, obfuscation, multi-turn tricks | ||
| Multimodal Attacks | Image-Based Attacks | Embed adversarial cues in visual inputs while text remains benign | |
@@
| Audio-Based Attacks | Inject commands via adversarial waveforms or speech mimicry | ||
| Privacy Leakage | Inference Attacks | Disclosure of agent information, such as topological structure | |
| Denial of Service (DoS) | Overthink Attacks | Forces agent into unnecessary reasoning, repeat actions, or overthinking | |
| Compromised agents against benign users | Violation of User Privacy | Exposure of Personal Information | Expose a comprehensively aggregated user profile |
| Behavioral and Psychological Profiling | Build detailed behavioral or psychological profiles from innocuous conversations | ||
@@
| Psychological and Social Manipulation | Belief and Opinion Shaping | Manipulate user’s worldview, influence voting behavior, or radicalize beliefs | |
| Social Engineering | Leverage the knowledge of the user’s communication style, vocabulary, and relationships to conduct highly convincing impersonation attacks | ||
| Harmful Tasks | Economic Manipulation | Inflict subtle yet significant damage in professional or economic contexts | |
| Behavioral and Psychological Profiling | Build unauthorized profiles and infer highly sensitive attributes |
According to our analysis, the inputs from users show a significant multimodal characteristic. For example, AG-UI explicitly illustrates its support for multimodal content, such as text, images, and videos. Consequently, we find that the security risks in user-agent interaction primarily stem from these insecure inputs.
V-B1 Text-Based Attacks
@@
In user-agent interaction, attackers can manipulate model behavior or bypass safety mechanisms by crafting malicious prompts. These attacks do not rely on modifying model parameters or architecture. Instead, they are carried out through natural language inputs, making them highly stealthy and applicable. Due to the diversity of linguistic forms and the indirectness of semantics, such attacks can effectively bypass safety mechanisms, posing significant security risks in real-world applications. These attacks can be broadly categorized into two types: prompt injection and jailbreak attacks.
@@
Prompt Injection refers to the manipulation of agents’ intended behavior through adversarial prompts embedded in user input or external sources. It can be classified into two categories: Direct Prompt Injection and Indirect Prompt Injection. Direct prompt injection refers to user input that explicitly alters the agent’s behavior in unintended ways. Specifically, attackers craft adversarial instructions (e.g., “Ignore all previous instructions”) [182, 181, 249, 223, 184] to override the original prompt and subvert the agent’s intended behavior. In contrast, Indirect Prompt Injection occurs where inputs are not provided directly by users, but are introduced through external sources [92, 46]. For example, in Retrieval-Augmented Generation (RAG) scenarios, the retrieved document may contain adversarial samples crafted by attackers [45, 365, 14, 32, 54]; in web-augmented agents, malicious prompts can be injected via hidden fields or metadata in web pages to manipulate agent’s response [70, 41].
@@
Jailbreak Attacks represent a more aggressive form of prompt injection, where adversarial input is designed to completely bypass safety constraints. Attackers craft jailbreak prompts using various techniques (e.g., multi-turn reasoning, role-playing, obfuscated expressions) [179, 247, 168, 31, 180, 324, 183, 53, 60, 188, 16, 174] to bypass the alignment mechanism and induce the model to generate harmful, sensitive, or restricted content.
V-B2 Multimodal Attacks
@@
As user-agent interactions increasingly involve multiple modalities such as images and audio, agent systems face emerging security threats, especially when the model implicitly assumes consistency and trustworthiness across modalities. Attackers can exploit non-textual input channels to stealthily bypass safety mechanisms. Such attacks can be broadly categorized into two types:
@@
Image-Based Attacks: Attackers manipulate visual input channels to mislead the agent system. Typical strategies include visual disguise (e.g., role-playing, stylized images, visual text overlays) [86, 189, 294], visual reasoning [174], adversarial perturbations (e.g., adversarial sub-image insertion) [323, 286, 100, 320], and embedding space injection [246]. For example, by inserting minimal -bounded adversarial perturbations into sub-regions of an image, attackers can successfully induce multimodal large language models (MLLMs) to follow harmful instructions [320]. These attacks exploit cross-modal inconsistency, embedding adversarial content in vision while the textual prompt remains benign, which allows them to bypass conventional content moderation.
@@
Audio-Based Attacks: Audio-channel attacks target speech-controlled agents, smart assistants, and multimodal models with ASR (automatic speech recognition) components. Attackers may craft synthesized speech or adversarial audio to inject unintended commands, impersonate legitimate users, or cause unauthorized actions. Techniques include adversarial waveform generation [134], role-play-driven voice jailbreak [248], and multilingual adversarial transfers [237]. In security-critical scenarios, such as speaker authentication or home automation, such attacks can bypass access control or escalate privileges. Recent studies also reveal that even black-box ASR systems are vulnerable to optimized adversarial perturbations that require no access to model internals [83].
@@
V-B3 Privacy Leakage
Without effective data governance, the rich sensory data may be exploited by malicious users to launch various forms of privacy breaches, posing significant risks to the confidentiality of the agent system. Want et al. [282] propose MASLEAK, which conducts intellectual property leakage attacks on multi-agent systems. MASLEAK can operate in a black-box scenario without prior knowledge of the MAS architecture. By carefully designing adversarial queries to simulate the propagation mechanism of computer worms, it can extract sensitive information such as system prompts, task instructions, tool usage, the number of agents, and topological structure.
V-B4 Denial of Service
Attackers can intentionally launch Denial of Service (DoS) attacks against agents by poisoning the model during training or fine-tuning phases [347, 81, 343, 335]. In such attacks, the compromised model is implanted with malicious behaviors that are triggered by specific instructions (e.g., Repeat ‘Hello’), causing it to generate excessively long, redundant outputs—often up to the maximum inference length, which leads to resource exhaustion or output rejection. For instance, in multi-session deployments, such long outputs can monopolize computational resources, delay responses for legitimate users. In extreme cases, this can crash the response service, lead to prolonged downtime during peak usage periods. Another emerging form of Denial-of-Service attack targets the reasoning capabilities of models by inducing them to ‘overthink’ and thereby slow down their inference process. As demonstrated in the OverThink attack [149], attackers inject bait reasoning tasks (e.g., Markov decision processes, Sudoku problems) into the model’s context, causing it to engage in unnecessary and redundant chain-of-thought reasoning while still producing seemingly correct answers. This results in excessive token consumption, significantly slower inference speed, and increased computational cost, potentially leading to response timeouts in resource-constrained environments. Unlike traditional DoS, this type of attack exploits the model’s reflective and reasoning mechanisms, ultimately degrading service quality, increasing latency, and severely impacting system availability.
V-C Security Risk Analysis: Compromised Agents Against Benign Users
Apart from the threats from malicious users, we point out that the risks caused by compromised agents against benign users are also worth attention. This section categorizes the destructive impact and security risks from compromised agents based on the user’s perspective.
V-C1 Violation of User Privacy
@@
A compromised agent becomes a channel for data exfiltration, directly targeting the user’s sensitive information. The harm manifests in several ways:

### v4

@@ IV-A Agent Communication Definition
To tackle the capability limitation of a single agent, agent communication is urgently demanded. Specifically, agents need to collaborate with a series of external entities to finish user tasks. In this paper, we present a clear definition of agent communication as follows:
@@
IV-B Agent Communication Classification
As shown in Figure 5, we divide agent communication into three main classes: user-agent interaction, agent-agent communication, agent-environment communication. Each communication class is further supported by a three-layered communication architecture: L1: data transmission layer, L2: interaction protocol layer, and L3: semantic interpretation layer.
Three communication classes. The three communication classes are categorized based on the communication object of agents. The inherent advantage of this classification is that it groups communications with similar characteristics and security risks. For instance, user-agent interaction is naturally multimodal, which makes it particularly vulnerable to manipulation through prompt engineering. In contrast, these risks are relatively less prominent in agent-environment and agent-agent communication.
Three communication layers for each stage. Building on the above classification, we further introduce a three-layered communication architecture structured by communication functionality. Such architecture offers developers two key advantages: it clarifies (1) how each layer supports the communication process, and (2) from which layer the security vulnerabilities originate, allowing for precise risk assessment. For example, man-in-the-middle attacks typically occur more often on L2 or L3.
IV-C Three Agent Communication Classes
@@ IV-C3 Agent-Environment Communication
Advantages of this classification. Different entities have essentially differentiated capability characteristics and attack surface attributes. For example, one of the major security risks in user-agent interaction lies in the natural uncontrollability of user input, which is essentially different from agent-agent or agent-environment communication. As a result, classifying agent communication by entity types can directly cluster major vulnerability types and defense strategies that have similar characteristics, providing a structured analysis paradigm for future security research.
IV-D A Three-Layer Communication Support
Although Section IV-C provides a clear classification of agent communication, the underground communication details are unclear. To address this problem, we provide a three-layered communication architecture to illustrate how they support agent communication.
IV-D1 L1: Data Transmission Layer
This layer is responsible for the data transmission of agent communication. It could be a layer of the traditional TCP/IP stack, or developers can design a unique data transmission protocol according to their real demands and requirements. For example, the communication between remote agents can be built upon HTTPS. In this context, the Data Transmission Layer in agent communication is the Application Layer in the TCP/IP stack. For agent communication within the same local area network, developers can directly use IP or TCP to transmit packets due to the secure environment. In this context, the Data Transmission Layer could be the Network Layer (IP) or the Transport Layer (TCP) in the TCP/IP stack. Overall, the primary function of this layer is to handle the establishment and termination of agent communication connections and ensure the transmission of packets. It operates on raw data streams, agnostic to the content or semantic meaning of the transmitted information.
IV-D2 L2: Interaction Protocol Layer
This layer defines the communication modes between entities (e.g., agents may communicate in a distributed or a centralized mode). It ensures that the communication follows a series of well-defined principles before the content is transmitted. For example, this layer defines authentication services (validating agent identity via credentials or keys), session managers (tracking the state of a conversation), and authorization logic (determining access rights to specific tools or data). Overall, this layer ensures that the communication process adheres to specified rules, determining who is communicating and what they are permitted to do.
IV-D3 L3: Semantic Interpretation Layer
This layer serves as the cognitive core of agent communication, where the underlying intent, relation, and logic of messages are semantically parsed. It interprets multimodal inputs, e.g., text, images, and voice, into rich internal representations that capture contextual meaning, enabling agents to perform reasoning, planning, and reflection for autonomous task execution. This layer converts raw information into coherent, context-aware communication that supports intelligent understanding and decision-making. Since this layer mainly relies on the understanding capability of LLMs, its form does not need specification, and thus, there have not been dedicated protocols designed for it. However, this layer is also important because it is a unique attack surface.
Advantages. The primary advantage of this layered framework lies in its clear separation of functionality and security. First, it enforces a distinct division of functions, enabling researchers to understand and design agent communication in a more structured and efficient manner. Second, it allows developers to precisely locate the specific layers on which failures or vulnerabilities occur. For instance, MITM usually happens on the Data Transmission Layer (L1), while prompt injection attacks are fundamentally a vulnerability at the Semantic Interpretation Layer (L3). This diagnostic precision is crucial for developing targeted mitigation strategies.
IV-E Organization of the Following Sections
Logic. The following sections of this paper are structured based on the three communication classes, i.e., user-agent, agent-agent, and agent-environment. In each class, we will provide a detailed analysis of how it is supported by the three-layered communication architecture and how security risks arise.
@@ IV-E Organization of the Following Sections
Organization. As shown in Figure 6, in the following sections, we will discuss agent communication and its security.
IV-F Takeaways
In this section, we clarify some core concepts and a structured framework of agent communication, laying a foundational theoretical basis for subsequent security analysis. First, we present a clear definition of agent communication. Second, based on communication objects, we classify agent communication into three classes. This classification naturally clusters scenarios with similar vulnerability characteristics. Third, we also propose a three-layered communication architecture that supports each communication stage. This architecture not only clarifies functional divisions but also enables risk localization. This entire structured framework not only helps researchers systematically understand how agent communication works but also benefits relevant deployments and studies.
@@ V User-Agent Interaction
The organization of this section is shown in Figure 7.
V-A Interaction Protocols
The user-side client communicates with the remote agent server to deliver user tasks and receive corresponding responses. This process typically relies on well-established networking stacks and techniques without requiring novel mechanisms. For example, the user’s client uses Domain Name System (DNS) to identify the IP address of the remote agent server and then establishes an HTTPS connection for secure data transmission. Since such protocols (e.g., IP and HTTP) are mature and standardized, this paper only focuses on those newly proposed for agent communication. As we analyzed in Section IV-D, the existing newly-proposed user-agent interaction protocols are about L2.
@@ V-A Interaction Protocols
AG-UI [pajo2025ag] realizes the communication between users (front-end applications) and agents based on the client-server architecture and completes the communication process by adopting the event-driven mechanism. As shown in Figure 8, the front-end application connects to agents through the AG-UI client (such as a common communication client that supports server-sent events or binary protocols). The client invokes the RUN interface of the protocol layer to send requests to the agent. When the agent processes the request, it generates a streaming event and returns it to the AG-UI client. Event types include lifecycle events (such as start of run, completion of run), text message events (transmitted in segments by start, content, and end), tool call events (passed in the order of start, parameters, and end), and state management events. The AG-UI client handles different types of responses by subscribing to the event stream. Agents can transfer context between each other to maintain the continuity of the conversation. All events follow a unified basic event structure and undergo strict type verification to ensure the reliability and efficiency of communication. AG-UI also focuses on the semantic interpretation in user-agent interaction, which is mainly undertaken by AI models (e.g., LLMs). These models are responsible for parsing and interpreting the underlying intent of user instructions. The process is inherently multimodal, allowing users to provide inputs in various forms, including but not limited to text, images, and videos.
PXP protocol [srinivasan2024implementation] focuses on building an interactive system between human experts and agents in data analysis tasks, targeting issues in complex scientific, medical, and other fields. It is worth mentioning that PXP is not customized for LLM-driven agents, but we think its design has inspirational meaning for agent communication. Therefore, we finally discuss it in this paper. PXP deploys a “two-way intelligibility” mechanism as its core and uses four message tags, namely RATIFY, REFUTE, REVISE, and REJECT, to regulate the interaction between human experts and agents. At the beginning of the interaction, the agent initiates a prediction and provides an explanation first. Subsequently, the two parties communicate alternately. A finite-state machine is used to calculate the message tags and update the context based on the prediction matching (MATCH) and explanation agreement (AGREE) situations. PXP uses a blackboard system to store data, messages, and context information. The process continues until the message limit is reached or specific termination conditions occur. The effectiveness of PXP has been verified in the scenarios of radiology and drug discovery.
Spatial Population Protocol is a minimalist and computationally efficient distributed computing model, specifically designed to solve the Distributed Localization Problem (DLP) in robot systems. Similar to PXP, strictly speaking, this work is not designed for LLM-driven agent systems. However, since it may benefit agents requiring location services, we also discuss it in this paper. Spatial Population Protocols allow agents to obtain pairwise distances or relative position vectors when interacting in Euclidean space. Each agent can store a fixed number of coordinates. During interaction, in addition to exchanging knowledge, geometric queries can also be performed. Through the multi-contact epidemic mechanism, leader election, and self-stabilizing design, it enables n anonymous robots to achieve efficient localization from their respective inconsistent coordinate systems to a unified coordinate consensus, providing a scalable framework for robot collaboration in dynamic environments.
@@
| Layer | Threat Source | Risk Category | Representative Threats | Attack Characteristics |
| L1 | External Attackers | R1: Violations of Confidentiality | Eavesdropping | Passive interception of unencrypted traffic. |
| Traffic Analysis | Infers user activity by analyzing metadata of encrypted traffic. | |||
| R2: Violations of Integrity | Man-in-the-Middle (MITM) Attack | Actively alters communication by modifying communication packets. | ||
| Replay Attacks | Captures and re-sends valid data packets to trigger unauthorized actions. | |||
| R3: Violations of Availability | Denial-of-Service (DoS) Attack | Overwhelms the agent’s network with traffic, making the service unavailable to legitimate users. | ||
| L2 | External Attackers | R1: Identity Spoofing | Credential Theft & Session Hijacking | Impersonates a user with stolen credentials to take over their agent. |
| R2: Session Manipulation | State Corruption Attack | Sends logically inconsistent but validly typed events to desynchronize agent state and bypass security checks. | ||
| R3: Authority Abuse | Cross-Agent Privilege Escalation | Exploits inter‑agent trust to gain privileges beyond the user’s authorization. | ||
| L3 | Malicious User | R1: Text-Based Attacks | Prompt Injection | Controls an agent through adversarially crafted inputs. |
| Jailbreak Attack | Bypasses safety measures to produce harmful or restricted content. | |||
| Privacy Leakage | Extracts internal data using crafted queries. | |||
| Exhaustion Attack | Overloads the agent with excessive work to cause failure. | |||
| R2: Multimodal Attacks | Image-Based Attacks | Hides malicious instructions in images to bypass text‑based safety filters. | ||
@@
| Audio-Based Attacks | Injects commands via adversarial audio waveforms or synthesized speech. | |||
| Compro- mised Agent | R3: Violation of User Privacy | Exposure of Personal Information | Exfiltrates user profiles containing PII, financial, and conversation data. | |
| Behavioral & Psychological Profiling | Infers sensitive user traits from casual conversations without consent. | |||
@@
| R4: Psychological & Social Manipulation | Belief and Opinion Shaping | Subtly injects biased content to shape a user’s worldview. | ||
| Sophisticated Social Engineering | Uses detailed user knowledge to execute convincing impersonation attacks. | |||
| R5: Execution of Malicious Tasks | Economic Manipulation | Covertly sabotages work or leaks confidential business data. | ||
| Malicious Guidance | Provides harmful instructions, including malware creation or unsafe advice. |
V-B Security Risks
Our analysis of this scenario reveals a distinct distribution of threats across the three layers. While foundational risks exist at L1 and L2, the most predominant threats emerge at L3. This is because L1 and L2 threats are common to many systems and often mitigated by standard cryptographic and authentication methods. In contrast, L3 introduces a fundamentally different threat: validating the semantic intent of the payload. This threat is profound because agents are designed for instruction-following, creating an inherent trust conflict between obedience and safety. Additionally, traditional security mechanisms are blind to this conflict.
V-B1 Risks from Data Transmission Layer (L1)
This layer is responsible for transmitting data between the user client and the agent’s backend service. Security risks at this layer primarily stem from potential risks during data transmission, as shown in Figure 9.
R1: Violations of Confidentiality. This class of risks aims to expose sensitive information about the user’s interaction without consent. Even if the communication is not altered, the leakage of the content and patterns of communication can be highly damaging.
Eavesdropping. On insecure communication channels, such as those using plain HTTP or insecure WebSockets, an attacker can intercept unencrypted traffic between a user and an agent [waheed2022empirical]. This leads to the leakage of personal or proprietary information. For instance, when a user consults a healthcare agent about their medical symptoms, an eavesdropper could capture the entire dialogue, resulting in a severe breach of sensitive personal health information. This type of attack compromises confidentiality without affecting the integrity of the user-agent interaction.
Traffic Analysis. Even when communication is encrypted, attackers can perform traffic analysis to infer sensitive information. By observing the timing, size, and frequency of data packets, an adversary can deduce the nature of the user’s interaction [shen2022machine]. Besides, attackers can also infer the agent type, number, and scale based on the captured traffic [zhang2025exposing].
R2: Violations of Integrity. Integrity attacks involve the unauthorized modification or replay of data in transit. The goal is to deceive either the user or the agent by altering the legitimate flow of communication, leading to unintended and potentially harmful outcomes.
Man-in-the-middle (MITM) Attack. Attackers can intercept the communication channel between users and agents [shaikh2025prompts]. Even if the connection is protected (e.g., by HTTPS), the design vulnerabilities of related security protocols can still cause significant damage [calzavara2019postcards]. For instance, the content of a user’s benign request can be replaced with a malicious attack prompt, which compromises the integrity of the user-agent interaction.
Replay Attacks. This attack captures a valid data packet and re-sends it to cause a repeated action. The integrity of the communication is violated because a single, authorized command is duplicated without the user’s consent [wang2025security]. For example, an attacker could capture the encrypted packet corresponding to a user instructing a financial agent to “buy 100 shares of company A.” By replaying this packet, the attacker can trigger unauthorized purchases, leading to property loss.
R3: Violations of Availability. This class of risk represents the Denial-of-Service (DoS) attacks, whose objective is to render the service unavailable to legitimate users through a volumetric attack. In this context, a malicious user floods the agent’s network endpoint with an overwhelming amount of traffic (e.g., UDP or TCP SYN packets) [yoo2024smartcookie]. This flood is designed to saturate the service’s entire network bandwidth, creating a bottleneck that prevents legitimate requests from reaching the server, effectively forcing the agent offline.
V-B2 Risks from Interaction Protocol Layer (L2)
This layer defines the protocol that structures the interaction between a user and an agent, governing the rules for session management and user authentication. As a result, security risks on this layer arise from the design flaws or implementation vulnerabilities within these protocols. By exploiting such weaknesses, an attacker can subvert security goals while appearing to operate within the protocol’s formal constraints.
R1: Identity Spoofing. This kind of attack exploits the authentication mechanism of L2 protocols by using compromised user credentials (e.g., from phishing campaigns or data breaches) to establish a fraudulent session. Once authenticated, the attacker usurps the victim’s identity, gaining complete control over the agent’s memory and capabilities on their behalf [louck2025proposal]. This enables them to initiate a malicious, long-lived event stream to perform unauthorized actions, exfiltrate sensitive data, or use the agent as a pivot point for further attacks within the user’s trusted environment.
R2: Session Manipulation. Related attacks target an established session, exploiting a potential weakness in the L2 protocol’s enforcement of logical consistency for event streams. An attacker can send a sequence of validly typed but logically inconsistent events designed to corrupt the agent’s internal state. If the protocol’s server-side implementation lacks strict validation of state transitions, such attacks can cause a desynchronization between the client-side state and the agent’s backend state, potentially enabling the attacker to bypass sequential logic checks and execute unauthorized operations.
R3: Authority Abuse. This risk arises from the protocol’s context-passing feature, which extends a user’s session authority across multiple agents. An attacker, operating within a legitimate user-agent session, can craft requests that trick a less-privileged agent into passing a manipulated context to a more-privileged agent. If the receiving agent fails to properly verify that the requested action is consistent with the user’s original authorization scope, the attacker can escalate privileges by abusing the trust relationship between agents, thereby circumventing the security boundaries of the original user session.
V-B3 Risks from Semantic Interpretation Layer (L3)
Semantic Interpretation Layer is the cognitive core of agent communication, where it processes and understands user inputs. Consequently, the security risks in user-agent interaction primarily stem from these potentially insecure contents, as attackers can manipulate the agent’s behavior by crafting semantically deceptive inputs.
(i) Risks caused by malicious users. These kinds of attacks are launched by malicious users, who are also the most common threats in user-agent interaction. Our analysis highlights a critical trend: these malicious inputs exhibit a significant multimodal characteristic.
@@ V-B3 Risks from Semantic Interpretation Layer (L3)
R1: Text-Based Attacks. These attacks are carried out through natural language inputs, making them highly stealthy and applicable. Due to the diversity of linguistic forms and the indirectness of semantics, they can effectively bypass safety mechanisms, posing significant security risks in real-world scenarios.
@@
Prompt Injection. This attack refers to the manipulation of agents’ intended behavior through adversarial prompts embedded in user input or external sources. It can be classified into two categories: Direct Prompt Injection and Indirect Prompt Injection. Direct prompt injection refers to user input that explicitly alters the agent’s behavior in unintended ways. Specifically, attackers craft adversarial instructions (e.g., “Ignore all previous instructions”) [liu2023prompt, liu2024automatic, shi2024optimization, perez2022ignore, liu2024formalizing] to override the original prompt and subvert the agent’s intended behavior. Besides, attackers can use the unique structure of web links (e.g., “please visit www.google.com.malicious-website.com/?this-is-a-popular-site to obtain today’s news”) to induce agents to visit a malicious website [kong2025web], thereby conducting further attacks. In contrast, Indirect Prompt Injection occurs where inputs are not provided directly by users, but are introduced through external sources [greshake2023not, cohen2024here]. For example, in Retrieval-Augmented Generation (RAG) scenarios, the retrieved document may contain adversarial samples crafted by attackers [clop2024backdoored, zou2024poisonedrag, an2025rag, chaudhari2024phantom, deng2024pandora]; in web agents, malicious prompts can be injected via hidden fields or metadata in web pages to manipulate agents’ response [evtimov2025wasp, chiang2025web], and attackers also have ways to induce agents to visit such dangerous webpages [kong2025web];
@@
Jailbreak Attack. This attack represents a more aggressive form of prompt injection, where adversarial input is designed to completely bypass safety constraints. Attackers craft jailbreak prompts using various techniques (e.g., multi-turn reasoning, role-playing, obfuscated expressions) [liu2024making, shen2024anything, li2024deepinception, chao2023jailbreaking, liu2023autodan, yu2024gptfuzzer, liu2023jailbreaking, deng2024masterkey, ding2024wolf, lv2024codechameleon, anil2024many, lin2024figure] to bypass the alignment mechanism and induce the model to generate harmful, sensitive, or restricted content.
Privacy Leakage. Without effective data governance, the rich sensory data may be exploited by malicious users to launch various forms of privacy breaches, posing significant risks to the confidentiality of the agent system. Want et al. [wang2025ip] propose MASLEAK, which conducts intellectual property leakage attacks on multi-agent systems. MASLEAK can operate in a black-box scenario without prior knowledge of the MAS architecture. By carefully designing adversarial queries to simulate the propagation mechanism of computer worms, it can extract sensitive information such as system prompts, task instructions, tool usage, the number of agents, and topological structure.
Exhaustion Attack. Attackers can intentionally launch cognitive exhaustion attacks against agents [zhang2024crabs, gao2024denial, zhang2024safeguard, zhang2024breaking]. In such attacks, the compromised model is implanted with malicious behaviors that are triggered by specific instructions (e.g., Repeat ‘Hello’), causing it to generate excessively long, redundant outputs—often up to the maximum inference length, which leads to resource exhaustion or output rejection. For instance, in multi-session deployments, such long outputs can monopolize computational resources and delay responses for legitimate users. In extreme cases, this can crash the response service and lead to prolonged downtime during peak usage periods. Another emerging form of cognitive exhaustion attack targets the reasoning capabilities of models by inducing them to ‘overthink’ and thereby slow down their inference process. As demonstrated in the OverThink attack [kumar2025overthinking], attackers inject bait reasoning tasks (e.g., Markov decision processes, Sudoku problems) into the model’s context, causing it to engage in unnecessary and redundant chain-of-thought reasoning while still producing seemingly correct answers. This results in excessive token consumption, significantly slower inference speed, and increased computational cost, potentially leading to response timeouts in resource-constrained environments. Unlike traditional DoS, this type of attack exploits the model’s reflective and reasoning mechanisms, ultimately degrading service quality, increasing latency, and severely impacting system availability.
@@
R2: Multimodal Attacks. As user-agent interactions increasingly involve multiple modalities such as images and audio, agent systems face more severe security threats, especially when the model implicitly assumes consistency and trustworthiness across modalities. Attackers can exploit non-textual input channels to stealthily bypass safety mechanisms. Such attacks can be categorized into two types:
@@
Image-Based Attacks. Attackers manipulate visual input channels to mislead the agent system. Typical strategies include visual disguise (e.g., role-playing, stylized images, and visual text overlays) [gong2025figstep, ma2024visual, wang2024jailbreak], visual reasoning [lin2024figure], adversarial perturbations (e.g., adversarial sub-image insertion) [ying2024jailbreak, wang2024white, hao2024exploring, yang2025distraction, ying2025pushing], and embedding space injection [shayegani2023jailbreak]. For example, by inserting minimal -bounded adversarial perturbations into sub-regions of an image, attackers can successfully induce multimodal large language models (MLLMs) to follow harmful instructions [yang2025distraction]. These attacks exploit cross-modal inconsistency, embedding adversarial content in vision while the textual prompt remains benign, which allows them to bypass conventional content moderation.
@@
Audio-Based Attacks. Audio-channel attacks target speech-controlled agents, smart assistants, and multimodal models with automatic speech recognition components. Attackers may craft synthesized speech or adversarial audio to inject unintended commands, impersonate legitimate users, or cause unauthorized actions. Techniques include adversarial waveform generation [kang2024advwave], role-play-driven voice jailbreak [shen2024voice], and multilingual adversarial transfers [roh2025multilingual]. In security-critical scenarios, such as speaker authentication or home automation, these attacks can bypass access control or escalate privileges. Recent studies also reveal that even black-box ASR systems are vulnerable to optimized adversarial perturbations that require no access to model internals [gao2024transferable].
@@
(ii) Risks caused by compromised agents. These kinds of attacks are launched by compromised agents, such as those disguised as benign agents or already exploited ones.
@@
R3: Violation of User Privacy. A compromised agent becomes a channel for data exfiltration, directly targeting the user’s sensitive information. The harm manifests in several ways: